[Hallc_running] [New Logentry] Hall C RC update

hszumila at jlab.org hszumila at jlab.org
Thu Nov 4 09:15:01 EDT 2021

Logentry Text:
Last night we established beam just before midnight and continued production at 42uA. We obtained an additional 96 pions at this setting with a current total of 9696 pions. 

We lost the beam in the night for an RF separation issue. It was resumed until around 5am when we had unstable beam followed by an LCW breaker trip in the SA causing a power outage that is currently preventing beam. We hope to resume data-taking in the next few hours once this event is recovered. 

Beam will be taken away sometime tomorrow morning around 7am. We will be swept for restricted access for the aerogel change (and accelerator change) and for the SHMS to be moved away from the beam line. 

We will have an RC meeting today at 10am.

Bluejeans link: https://bluejeans.com/861439920


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3936831
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20211104/7ca0dd7d/attachment.html>

More information about the Hallc_running mailing list